A spherical bubble rises from the bottom of a lake whose temperature is 10 oC at the bottom and 22 oC at the surface. If the bubble doubles its volume by the time it reaches the surface, how deep is the lake? (please be specific on how you calculate the volume.)
Answer #1


T1 =10 oC = 10+273 = 283 K

T2 = 22 oC = 22+273 = 295 K

V1 = 2V2

P1 = 1atm = 1.01*10^5 N/m^2

From ideal gas equation

PV = nRT

P2V2/P1v1 = T2/T1

P2/(2*1.01*10^5) = 295/283

P2 = 2.11*10^5 N/m^2

density of water d =1000 kg/m^3

P2 -P1 = dgh

(2.11-1.01)*10^5 = 1000*9.8*h

h =11.22 m